OpenStudy (anonymous):

In the 1950s, an experimental train, which had a mass of 2.50X10^4 kg, was powered across a level track by a jet engine that produced a thrust of 5.00X10^5 N for a distance of 509m. a. Find the work done on the train b. Find the change in Kinetic energy. c. Find the final kinetic energy of the train if it started from rest. d. Find the final speed of the train if there had been no friction.

OpenStudy (loser66):

a) W = F *d = 5.00x10^5 * 509= 2.545 x10^8 J c) a = F/m = 5.00 x 10^5 / 2.50 x 10^4 = 20m/s^2 and\(V^2 = V_0^2 +2ad \) so \( V = \sqrt {2ad}=\sqrt{2*20*509}= 142.69 m/s\) Kinetic Energy \(KE = \dfrac{1}{2}mV^2 = \dfrac{2.50 *10^4 *2*20*509}{2}= 2.54 * 10^8joules\)